# Computer Validation of Open Gaps for the Almost Mathieu Operator with Critical Coupling

### v1

@@ Abstract.
We present some computer assisted methods to prove the existence of spectral gaps for the Almost Mathieu operator at critical coupling and give rigorous numerical estimates on their size. As an example we show that the first 8 gaps predicted by the Gap Labelling theorem are open when $\omega=(\sqrt{5}-1)/2$ and 12 of them are open when $\omega=e-2$. A dynamical method based on the constructive conjugation to a hyperbolic cocycle and a spectral method based on the rigorous computation of the eigenvalues of finite-dimensional matrices are presented. We also present some experiments and conjectures on gap size for the associated peridodic problems.
@@ 1. Introduction
Figure 1. Hofstadter butterfly: frequency on the vertical direction, spectrum on the horizontal direction
@@ 2.1.1. Quasi-periodic cocycles and their reducibility
In the validated numerical algorithms that we will present, it is important to describe of the dynamics of the eigenvalue equation (2.3), or rather its equivalent 1st order system,
@@ 2.1.2. The rotation number and the Gap Labelling Theorem
[…]
@@ 2.1.3. Dynamics inside a spectral gap
It is therefore possible to establish the existence of an open spectral gap in a Schrödinger operator (e.g the AMO) provided the cocycle is shown to be uniformly hyperbolic for a particular energy $a$. It is well-known that uniform hyperbolicity is equivalent to the conjugacy to a projective diagonal cocycle in $PSL(2,\mathbb{R})$ (see Theorem 3.8.2 in [12] and references therein). In order to prove conjugacy to a hyperbolic cocycle we will need a sufficient condition for uniform hyperbolicity [22, 16, 21] which can be validated using rigorous numerics and is enough for our purposes.
@@
If the invariant stable and unstable bundles have non-orientable topology, we need to apply the same theorem but on the cocycle $\left(\dfrac{\omega}{2},A_{a}(2\cdot)\right)$on $\mathbb{R}/(2\mathbb{Z})\times\mathbb{R}^{2}$ instead of /×$\mathbb{R}/\mathbb{Z}\times\mathbb{R}^{2}$.
@@ 2.3. Spectral Method: Periodic approximants
Besides using the dynamical method, we also compute some spectral gaps for Almost Mathieu operators using periodic approximations, which is what we call the spectral method. This is essentially a computer-assisted validation of the the original periodic approximation which Hofstadter [26] used to produce his famous butterfly-shaped plots.
@@
and the constant $60$ can be replaced by a better bound which depends on the difference $\omega-\omega^{\prime}$. The following result summarizes the corresponding bounds that the above estimates imply for the gaps similarly to Proposition 2.2 did for the estimate 2.6
@@
then the gap with label $k$ is open for $\omega^{\prime}$ it includes the interval
@@ 2.3.2. Almost Mathieu operators with rational frequencies
Spectral bands are separated by $q-1$ (possibly void) spectral gaps
@@
and, for $l\equiv 1,2$ (mod 4), these correspond to the eigenvalues of the matrix $H^{-}_{p/q}(\theta)$ with anti-periodic boundary condition:
@@
is either $+2$ for periodic eigenvalues or $-2$ for anti-periodic eigenvalues.
@@
Consider an Almost Mathieu operator with coupling $b=2$, rational frequency $\omega=p/q$ and some $\theta\in\mathbb{T}$. Let $\Delta_{k}(\theta)$ be the spectral gaps of the spectrum $\Sigma_{\omega}(\theta)$ for $k=0,\dots,q-1$
  1. (1)
    If $q$ is even or $q$ is odd and $b>0$, the gaps satisfy
@@
  2. (2)
    If $q$ is odd and $b<0$, then
    $\Delta_{k}\left(\frac{1}{2q}\right)=\bigcap_{\theta}\Delta_{k}(\theta),\quad k=1,3,\ldots$
    $\Delta_{k}(0)=\bigcap_{\theta}\Delta_{k}(\theta),\quad k=2,4,\ldots$
@@ 2.3.3. Computer-assisted proof of the persistence of gaps for irrational frequencies
We take a certain rational approximation $p/q$ of $\omega$, with $p$ and $q$ coprime, and compute a validated approximation of the eigenvalues of the $q$-dimensional matrices $H^{+}_{p/q}(\theta)$ and $H^{-}_{p/q}(\theta)$ for $\theta=0$ and $\theta=1/2q$. In our case we used the Arblib library [29], a C library for interval arithmetic, which provides the verified computation of eigenvalues for a matrix in interval arithmetic, see [46, 49]. We chose to compute only rational frequencies $\omega=p/q$ with $q$ being an odd number to avoid the unnecessary complication of having 00 as a double eigenvalue (see Figure 1).
@@
We can use the bounds given by Proposition 2.2 or 2.3 to give, for every gap in the periodic problem with rational frequency $p/q$ that we have taken, a rigorous bound for which this gap will exist for any $\omega$ close enough to $p/q$ as well a rigorous lower bound on its size. Indeed, taking an irrational $\omega$ (like $(\sqrt{5}-1)/2$ or $e-2$) and some rational approximation $p/q$, we can prove the existence of all the the gaps for irrational $\omega$ which are larger than a certain bound (given by the above two continuity results on the spectrum) and depends only on the distance between $\omega$ and this rational approximation $p/q$.
@@ 3.1.1. Results for $\omega_{1}=(\sqrt{5}-1)/2$
Using the dynamical method we obtained the gaps displayed in Table 1. A total of 8 gaps were obtained. The distance between the upper and the lower point of the spectrum minus the total sum of validated gaps for $\omega_{1}$ was $0.635$ which gives an upper bound of the measure of the spectrum (which is known to be zero) and represents the $86.63\%$ of the interval between the validated right endpoint of the lowest gap and the right endpoint of the upper gap.
@@ 3.1.2. Results for $\omega_{2}=e-2$
With the same dynamical method we obtained the gaps displayed in Table 2. A total of 8 gaps were obtained. The distance between the upper and the lower point of the spectrum minus the total sum of validated gaps for $\omega_{2}$ was $0.8273$ which gives an upper bound of the measure of the spectrum (which is known to be zero) and represents the $84.3\%$ of of the interval between the validated right endpoint of the lowest gap and the right endpoint of the upper gap.
@@ 3.2.1. Results for $\omega_{1}=(\sqrt{5}-1)/2$
We proved the existence of gaps for $\omega_{1}=(\sqrt{5}-1)/2$ which are given in Table 3. We used $p/q=987/1597$ and ball arithmetic with $50$ digits and showed the existence of $8$ spectral gaps (apart from bounds of the lowest and upper spectral gap). In this case, we used the estimates from Proposition 2.3 to obtain that gaps boundaries had to be enlarged by $0.0251248715580\pm 4.46\cdot 10^{-14}$ The measure of the spectrum for the rational approximation was $\Sigma(p/q)$ was 0.$0584216754614\pm 3.31\cdot 10^{-15}0.$. The distance between the upper and the lower point of the spectrum minus the total sum of validated gaps for $\omega_{1}$ was $0.844250239140\pm 2.13\cdot 10^{-13}$ which gives an upper bound of the measure of the spectrum (which is known to be zero) and represents the $83.17\pm 6.42\cdot 10^{-4}\%$ of the interval between the validated right endpoint of the lowest gap and the right endpoint of the upper gap.
@@ 3.2.2. Results for $\omega_{2}=e-2$
Using the same method, we also computed gaps in the resolvent set using interval arithmetic with $40$ digits and $p/q=719/1001$, which is a rational approximation of $\omega_{2}=e-2$. Using again the estimates from Proposition 2.3 we obtained an enlargement value of $0.0199157821095\pm 7.79\cdot 10^{-16}$ and this proved the existence of the $12$ gaps described in Table 4.
@@ 3.2.3. Remarks on the gaps of $\Sigma_{p/q}$
[…]
@@
We can observe that gap lengths seem to decay faster for $\omega_{2}$ than for $\omega_{1}$. In fact, we checked that for in both cases the following scaling law holds for all $k$ and
$\gamma_{k}>\frac{1}{|k|^{\nu}},\quad|k|>1$
@@
Lower bounds on the gap length are important in the proofs of gap opening, but they seem to depend very much on the arithmetical properties of the frequency and the label of each gap. In [11] (Theorem 3.3) it was shown that the gaps for the periodic problem are greater than $8^{-q}$ while in [3] (Theorem 7.2), the following was proved: for any irrational frequency and sequence of $p_{n}/q_{n}$ converging to it, there is a $\epsilon>0$ such that gaps of the periodic problem $p_{n}/q_{n}$ have size at least $e^{-\epsilon q_{n}}$. For the golden mean $\omega_{1}$ and $p/q=987/1597$ we checked that all gaps are greater than $K_{1}^{-q}$ for $K_{1}=1.01090201763\pm 3.96\cdot 10^{-12}$. while for $\omega_{2}=e-2$ and $719/1001$ this value was $K_{2}=1.02993096753\pm 8.70\cdot 10^{-13}$.
@@ 3.2.4. Exploration of different rational frequencies
Figure 4 shows the decay in the minimal gap length for every $q$ in this range. We can observe that the values $p=1,q-1$ and $p=2,q-2$ are special and that give the upper and lower bounds on gap length for every odd $q$. The following statement has been checked to be true for $q\leq 95$:
@@
Figure 5 shows the Lebesgue measure of the periodic spectrum as a function of the rational approximation $p/q$, which is known to tend to zero when $p/q$ tends to an irrational frequency $\omega$ The rate at which this limit tends to zero is related to a conjecture by Thouless [47] namely, for any sequence of coprime $p_{n}$ and $q_{n}$ such that $p_{n}/q_{n}\to\omega$ irrational one has
@@
Figure 5. Measure of the AMO for rational frequencies as a function of the rational frequency $p/q$ for odd $q$ and $q\leq 95$ (top) and the same with logarithmic y scale. Greylines connect dots whose frequency $p/q$ have the same $p$. It ressembles of Thomae’s function, although the vertical values, the measures $\Sigma_{p,q}$ are not exactly the same for all values $p/q$ with the same $q$.
@@
If such a an inequality was shown to be true for every coprime $p$ and $q$ (which cannot be settled with the methods in the present paper), the analysis of this conjecture would be reduced to the consideration of the cases $p=1$ and $p=q-1$ [24]. See also the discussion in [38].
Figure 6. Top: Difference $32C/\pi-q|\Sigma_{p/q}|$ for the as a function of $q$. Note that the slowest convergence is given exactly when $p=1$ or $p=q-1$, which seems to be faster than $\sqrt{2/q}$. Bottom: Same plot as a function of the rational approximations $p/q$.
Finally we wanted to investigate the Hölder continuity of the spectrum in Equation (2.9) and obtain lower bounds given the exact computation of the spectrum for some rational $p/q$. We used the fact that or even $q$, the spectrum of $\Sigma_{p/q}$ contains 00, which is not in the spectrum when $q$ is odd. Figure 7 displays, for every $p/q$ with $q$ the minimum positive value in the spectrum, which is precisely $d(0,\Sigma_{p/q})$ and allows to compute the constant in 2.9, proved to be lower that $60$ in [34]. For instance, computing the distance from 00, which lies in the spectrum of $\Sigma_{1/2}$, with the rest of the spectrum we obtained that the constant needs to be larger than $7.03127155516\pm 3.80\cdot 10^{-12}$
Figure 7. Plot of the minimum positive value in the spectrum $\Sigma_{p/q}$ for every odd $q\leq 95$ as a function of the frequency $p/q$. Values in red correspond to points of the form $(j/k,0)$ with even $k$ and $gcd(j,k)=1$ which are a collapsed gap for $\Sigma_{j,k}$.
@@ 4. Discussion and conclusions
In this particular model, the Almost Mathieu operator, the dynamical and the spectral method seem to yield similar results in terms of performance. For more general models, we expect the dynamical method to perform better since Chambers formula, which reduces the computation of the periodic problem to simply two phases, $\theta_{0}=0$ or $\theta_{0}=1/2q$ is specific of the Almost Mathieu operator. The dynamical method can easily adapted to other sampling functions than the cosine. Extension to higher frequencies, moreover, is straightforward, although technically challenging by the implementation problems of higher-dimensional Fourier series.
Although, to our knowledge, the opening of some specific gaps for the Almost Mathieu operator had not previously been shown, the results in the present paper give little clue on the the opening of all gaps. Our main motivation in presenting them was twofold. On one hand, to show that it is possible to prove realistic bounds on the gaps of the Hofstadter Butterfly which account for most of what can be experimentally observed, both from a numerical point of view [44] and from physical experiments [45]. On the other, by being able to rigorously bound some gaps, one can test some of the inequalities and bounds on gaps, most notably the opening of all gaps, for some finite bound. In this paper, we have tested some of these conjectures like Thouless conjecture, the possible ordering of gaps according to their label (which we proved to be false in general) and the same could be done with analytic upper bounds for the spectrum such as in [10].

### v2

@@ Abstract.
We present some computer assisted methods to prove the existence of spectral gaps for the Almost Mathieu operator at critical coupling and give rigorous numerical estimates on their size. As an example we show that the first 8 gaps predicted by the Gap Labelling theorem are open when the frequency $\omega$ is $\sqrt{5}-1)/2$ and 12 of them are open when $\omega=e-2$. A dynamical method based on the constructive conjugation to a hyperbolic cocycle and a spectral method based on the rigorous computation of the eigenvalues of finite-dimensional matrices are presented. We also present some experiments and conjectures on gap size for the associated periodic problems.
@@ 1. Introduction
Figure 1. Hofstadter butterfly: frequency on the vertical direction, spectrum on the horizontal direction.
@@ 2.1.1. Quasi-periodic cocycles and their reducibility
In the validated numerical algorithms that we will present it is important to describe the dynamics of the eigenvalue equation (2.3) or rather, its equivalent 1st order system,
@@ 2.1.2. The rotation number and the Gap Labelling Theorem
Moreover, the open intervals in the resolvent set of the spectrum, the spectral gaps where the rotation number is constant, are uniquely labelled by an integer called its label. This is the content of the Gap Labelling Theorem [31], which asserts that inside every open gap in the resolvent set, there is a unique label $k\in\mathbb{Z}$ such that the rotation number is resonant with respect to $\omega$, that is, it is of the form $\{k\omega\}\in[0,1)$ where $\{\cdot\}$ denotes the fractional part of a real number (see Theorem 4.10.3 in [12]) and the unique value $k$ is called the label of the gap.
@@ 2.1.3. Dynamics inside a spectral gap
It is therefore possible to establish the existence of an open spectral gap in a Schrödinger operator (e.g. the AMO) provided the cocycle is shown to be uniformly hyperbolic for a particular energy $a$. It is well-known that uniform hyperbolicity is equivalent to the conjugacy to a projective diagonal cocycle in $PSL(2,\mathbb{R})$ (see Theorem 3.8.2 in [12] and references therein). In order to prove conjugacy to a hyperbolic cocycle we will need a sufficient condition for uniform hyperbolicity [22, 16, 21] which can be validated using rigorous numerics and is enough for our purposes.
@@
If the invariant stable and unstable bundles have non-orientable topology, we need to apply the same theorem but on the cocycle $\left(\dfrac{\omega}{2},A_{a}(2\cdot)\right)$on $\mathbb{R}/(2\mathbb{Z})\times\mathbb{R}^{2}$ instead of $\mathbb{R}/\mathbb{Z}\times\mathbb{R}^{2}$.
@@ 2.3. Spectral Method: Periodic approximants
Besides using the dynamical method, we also compute some spectral gaps for Almost Mathieu operators using periodic approximations, which is what we call the spectral method. This is essentially a computer-assisted validation of the original periodic approximation which Hofstadter [26] used to produce his famous butterfly-shaped plots.
@@
and the constant $60$ can be replaced by a better bound which depends on the difference $\omega-\omega^{\prime}$. Similarly to Proposition 2.2 we can now summarize the implications that these bounds have on the estimates of gap size:
@@ 2.3.2. Almost Mathieu operators with rational frequencies
Spectral bands are separated by q1q-1 (possibly void) spectral gaps
@@
and, for $k\equiv 1,2$ (mod 4), these correspond to the eigenvalues of the matrix $H^{-}_{p/q}(\theta)$ with anti-periodic boundary condition:
@@
is either +2+2 for periodic eigenvalues or 2-2 for anti-periodic eigenvalues.
@@
Consider an Almost Mathieu operator with coupling $|b|=2$, rational frequency $\omega=p/q$ and some $\theta\in\mathbb{T}$. Let $\Delta_{k}(\theta)$ be the spectral gaps of the spectrum $\Sigma_{\omega}(\theta)$ for $k=0,\dots,q-1$. The gaps satisfy
@@ 2.3.3. Computer-assisted proof of the persistence of gaps for irrational frequencies
We take a certain rational approximation $p/q$ of $\omega$, with $p$ and $q$ coprime, and compute a validated approximation of the eigenvalues of the $q$-dimensional matrices $H^{+}_{p/q}(\theta)$ and $H^{-}_{p/q}(\theta)$ for $\theta=0$ and $\theta=1/2q$. In our case we used the Arblib library [29], a C library for interval arithmetic, which provides the verified computation of eigenvalues for a matrix in interval arithmetic, see [46, 49]. We chose to compute only rational frequencies $\omega=p/q$ with $q$ being an odd number to avoid the unnecessary complication of having 0 as a double eigenvalue (see Figure 1).
@@
We can use the bounds given by Proposition 2.2 or 2.3 to give, for every gap in the periodic problem with rational frequency $p/q$ that we have taken, a rigorous bound for which this gap will exist for any $\omega$ close enough to $p/q$ as well as a rigorous lower bound on its size. Indeed, taking an irrational $\omega$ (like $(\sqrt{5}-1)/2$ or e2e-2) and some rational approximation $p/q$, we can prove the existence of all the gaps for irrational $\omega$ which are larger than a certain bound (given by the above two continuity results on the spectrum) and depends only on the distance between $\omega$ and this rational approximation $p/q$.
@@ 3.1.1. Results for $\omega_{1}=(\sqrt{5}-1)/2$
Using the dynamical method we obtained the gaps displayed in Table 1. A total of 8 gaps were obtained, a part from the two unbounded spectral gaps at the bottom and the top of the spectrum. The distance between the upper and the lower point of the spectrum minus the total sum of validated gaps for $\omega_{1}$ was $0.635$ which gives an upper bound of the measure of the spectrum (which is known to be zero) and represents the $86.63\%$ of the interval between the validated right endpoint of the lowest gap and the right endpoint of the upper gap. The computation took 4h00m at a i7 single processor @1.8GHz.
@@ 3.1.2. Results for $\omega_{2}=e-2$
With the same dynamical method we obtained the gaps displayed in Table 2. A total of 8 gaps were obtained. The distance between the upper and the lower point of the spectrum minus the total sum of validated gaps for $\omega_{2}$ was $0.8273$ which gives an upper bound of the measure of the spectrum (which is known to be zero) and represents the $84.3\%$ of of the interval between the validated right endpoint of the lowest gap and the right endpoint of the upper gap. The computation took 4h00m at a i7 single processor @1.8GHz.
@@ 3.2.1. Results for $\omega_{1}=(\sqrt{5}-1)/2$
We proved the existence of gaps for $\omega_{1}=(\sqrt{5}-1)/2$ which are given in Table 3. We used $p/q=987/1597$ and ball arithmetic with $50$ digits and showed the existence of $8$ spectral gaps (apart from bounds of the lowest and upper spectral gap). In this case, we used the estimates from Proposition 2.3 to obtain that gaps boundaries had to be enlarged by $0.0251248715580\pm 4.46\cdot 10^{-14}$ . The measure of the spectrum for the rational approximation $\Sigma(p/q)$ was $0.00584216754614\pm 3.31\cdot 10^{-15}$. The distance between the upper and the lower point of the spectrum minus the total sum of validated gaps for $\omega_{1}$ was $0.844250239140\pm 2.13\cdot 10^{-13}$ which gives an upper bound of the measure of the spectrum (which is known to be zero) and represents the $83.17\pm 6.42\cdot 10^{-4}\%$ of the interval between the validated right endpoint of the lowest gap and the right endpoint of the upper gap. The computation took 3h50m at a i7 single processor @1.8GHz.
@@ 3.2.2. Results for $\omega_{2}=e-2$
Using the same method, we also computed gaps in the resolvent set using interval arithmetic with $40$ digits and $p/q=719/1001$, which is a rational approximation of $\omega_{2}=e-2$. Using again the estimates from Proposition 2.3 we obtained an enlargement value of $0.0199157821095\pm 7.79\cdot 10^{-16}$ and this proved the existence of the $12$ gaps described in Table 4. We note that, contrary to the golden mean case, it was enough to consider less digits to obtain more gaps. The computation took 1h40m at a i7 single processor @1.8GHz.
@@ 3.2.3. Remarks on the gaps of $\Sigma_{p/q}$
Figure 2 displays the general decay of gap length according to the label of each gap. Gaps tend to decay according to the label in general terms, but this decay is not strict, as it can be observed from the oscillating behaviour seen in the figure and can be checked for lower values of $q$. For instance, for $p=13$ and $q=21$, the gap with label 5-5 has length $0.012767645482867\pm 9.83\cdot 10^{-16}$ and the gap with label 6-6 has length $0.02806352869759\pm 4.35\cdot 10^{-15}$.
@@
We can observe that gap lengths seem to decay faster for $\omega_{2}$ than for $\omega_{1}$. In fact, we checked that, in both cases, the following scaling law holds
$\gamma_{k}>\frac{1}{|k|^{\nu}},\quad|k|>1,$
@@
Lower bounds on the gap length are important in the proofs of gap opening, but they seem to depend very much on the arithmetical properties of the frequency and the label of each gap. In [11] (Theorem 3.3) it was shown that the gaps for the periodic problem are greater than $8^{-q}$ while in [3] (Theorem 7.2), the following was proved: for any irrational frequency and sequence of $p_{n}/q_{n}$ converging to it, there is a $\epsilon>0$ such that gaps of the periodic problem $p_{n}/q_{n}$ have size at least $e^{-\epsilon q_{n}}$. For the golden mean $\omega_{1}$ and $p/q=987/1597$ we checked that all gaps are greater than $K_{1}^{-q}$ for $K_{1}=1.01090201763\pm 3.96\cdot 10^{-12}$, while for $\omega_{2}=e-2$ and $719/1001$ this value was $K_{2}=1.02993096753\pm 8.70\cdot 10^{-13}$.
@@ 3.2.4. Exploration of different rational frequencies
Figure 4 shows the decay in the minimal gap length for every $q$ in this range. We can observe that the values p=1,q1p=1,q-1 and p=2,q2p=2,q-2 are special and that give the upper and lower bounds on gap length for every odd $q$. The following statement has been checked to be true for $q\leq 95$:
@@
Figure 5 shows the Lebesgue measure of the periodic spectrum as a function of the rational approximation $p/q$, which is known to tend to zero when $p/q$ tends to an irrational frequency $\omega$ . The rate at which this limit tends to zero is related to a conjecture by Thouless [47] namely, for any sequence of coprime $p_{n}$ and $q_{n}$ such that $p_{n}/q_{n}\to\omega$ irrational one has
@@
Figure 5. Measure of the AMO for rational frequencies as a function of the rational frequency $p/q$ for odd $q$ and $q\leq 95$ (top) and the same with logarithmic y scale. Greylines connect dots whose frequency $p/q$ have the same pp. It ressembles of Thomae’s function, although the vertical values, the measures $\Sigma_{p,q}$ are not exactly the same for all values $p/q$ with the same $q$.
@@
If such inequalities were shown to be true for every coprime $p$ and $q$ (which cannot be settled with the methods in the present paper), the analysis of this conjecture would be reduced to the consideration of the cases $p=1$ and p=q1p=q-1 [24]. See also the discussion in [38].
Figure 6. Top: Difference $32C/\pi-q|\Sigma_{p/q}|$ for the as a function of $q$. Note that the slowest convergence is given exactly when $p=1$ or p=q1p=q-1, which seems to be faster than $\sqrt{2/q}$. Bottom: Same plot as a function of the rational approximations $p/q$.
Finally we wanted to investigate the Hölder continuity of the spectrum in Equation (2.9) and obtain lower bounds given the exact computation of the spectrum for some rational $p/q$. We used the fact that, for even $q$, the spectrum of $\Sigma_{p/q}$ contains 0, which is not in the spectrum when $q$ is odd. Figure 7 displays, for every $p/q$ with $q$ the minimum positive value in the spectrum, which is precisely $d(0,\Sigma_{p/q})$ and allows to compute the constant in 2.9, proved to be lower that $60$ in [34]. For instance, computing the distance from 0, which lies in the spectrum of $\Sigma_{1/2}$, with the rest of the spectrum we obtained that the constant needs to be larger than $7.03127155516\pm 3.80\cdot 10^{-12}$.
Figure 7. Plot of the minimum positive value in the spectrum $\Sigma_{p/q}$ for every odd $q\leq 95$ as a function of the frequency $p/q$. Values in red correspond to points of the form $(j/k,0)$ with even $k$ and $gcd(j,k)=1$ which give a value of zero because $\{0\}$ is a collapsed gap for these frequencies with even $k$.
@@ 4. Discussion and conclusions
In this particular model, the Almost Mathieu operator, the spectral method seems to give slightly better results, although the overall performance, in terms of gaps and computing time, are similar. However, for more general models, we expect the dynamical method to perform better since Chambers formula, which reduces the computation of the periodic problem to simply two phases, $\theta_{0}=0$ or $\theta_{0}=1/2q$, is specific of the Almost Mathieu operator and the spectral method would be much slower. The dynamical method can easily adapted to other sampling functions than the cosine. Extension to higher frequencies, moreover, is straightforward, although technically challenging by the implementation problems of higher-dimensional Fourier series.
Although, to our knowledge, the opening of some specific gaps for the Almost Mathieu operator had not previously been shown, the results in the present paper give little clue on the opening of all gaps. Our main motivation in presenting them was twofold. On one hand, to show that it is possible to prove realistic bounds on the gaps of the Hofstadter Butterfly which account for most of what can be experimentally observed, both from a numerical point of view [44] and from physical experiments [45]. On the other, by being able to rigorously bound some gaps, one can test some of the inequalities and bounds on gaps, most notably the opening of all gaps, for some finite bound. In this paper, we have tested some of these conjectures like Thouless conjecture, the possible ordering of gaps according to their label (which we proved to be false in general) and the same could be done with analytic upper bounds for the spectrum such as in [10].